write the equation of the line containing the points (-1,1) and (4,-4)
Answer by mananth(16946)

x1 y1 x2 y2
-1 1 4 -4

slope m = (y2-y1)/(x2-x1)
( -4 - 1 )/( 4 - -1 )
( -5 / 5 )
m= -1
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point(-1 ,1) inY =mx+b
1.00 = -1 * -1 + b
1.00 = 1 + b
b= 1 -1
b= 0
So the equation will be
Y = -x
m.ananth@hotmail.ca